# Conditions for enhancement of chemical reactions in gas phase inside a dark cavity

### v1

@@ IV.1 Theory - the cavity effect on the rate of a reaction
Notice that both $\Gamma_{polariton}(\epsilon_{cav})$ and $\Gamma_{min}^{polariton}(\epsilon_{cav})$ are also functions of $\omega_{cav}$ as one can see from Fig. 11 which presents the rate of reaction inside the cavity as a function of the strength of the coupling of the activated complex $[AB]^{\#}$ with the cavity for two different values of $\omega_{cav}.$ It shows how the decay-rate of the hydrogen exchanged reaction with Methane increased inside the cavity when the distance between the two mirrors are tuned to the resonance condition. The coupling between the cavity and the system, in the acceleration gauge representation is used here, where the dipole transition is given by, $d_{{TS,DB}}(\omega_{cav})=\langle E_{TS}|\partial V_{ad}/\partial X|E_{DB}\rangle$ and the coupling strength parameter (in atomic unites) is given by $\alpha_{cav}=\epsilon_{cav}/\omega_{cav}^{2}$[26]. For the resonance condition we use one might be able to select several different cavity frequencies. $\omega_{cav}$ which couples the TS (transition state) resonance with a dynamical barrier (DB) broad resonance provided its position is lower than the position of the TS resonance (such that the TS in the vacuum of QED have about the same energy as the DB resonance that in the first excited QED mode of the cavity), Notice that the minimal distances between the two mirrors is given by, $L_{Z}=\pi c/\omega_{cav}\equiv\lambda_{cav}/2$. Therefore, the minimal distance depends on $\omega_{cav}$.
As expected by the analysis given in Eq.9, the enhancement of the reaction rate by the cavity is reduced as the coupling of the molecular state with the external field is increased. This is a spacial case of the generality of Bethe’s proof that there are situations where the increasing of the strength of the coupling of the system with external field not necessarily increase the rate of the reaction. See Fig.11 that shows the Bethe effect in the calculations of the effect of the cavity on the decay rate of the chemical reaction. That is, the increasing of the coupling strength parameter does not results necessarily in increasing the rate. When the difference between the resonance widths is sufficiently large (i.e., the TS is coupled with a broad DB resonance) the increasing of the strength of the coupling between the cavity and the molecules results in the suppression of the rate. As one can see from the results presented in Fig.11 the maximal enhancement of the rate of the reaction is obtained for a relative weak coupling between the cavity and the molecules.
@@
Using the adiabatic resonance eigenvalues and eigenfunctions for the $O+D_{2}$ reaction we have caculated and introdcued in the previous section we evaluate the effect of the cavity on the reaction rate using Eq.13.
Figure 11: The effect of the cavity on the reaction rate, $\Gamma_{polariton}$, of the TS polaritions of $O+D_{2}\xrightarrow[]{\Gamma_{polariton}(\epsilon_{cav})}OD+D$ reaction. Reaction rate in Hartree and cavity-coupling parameter, in atomic units. The cavity be tuned into resonance condition between the states marked in blue in Fig.10 by adjusting the distance between the two Fabrey-Perot mirrors (see Eq. 5). Rate in Hartree and coupling in atomic units.
@@ V The influence of a dark cavity on additional chemical reactions
An example for the first case case is scattering of hydrogen atom from the Van der Waals complex ArCl[30]. The poles which are presented in Fig.2 of Ref.17 are resonances which are associated with the temporary trapping of a hydrogen atom in between Ar and Cl. Only the first row of resonances in this figure have a classical analogue and are associated with the classical unstable periodic orbits. All other resonances are quantum diffraction resonances which do not have a classical analogue. These are Feshbach resonances of the two-dimensional PES where one coordinate is the reaction coordinate and the second one is perpendicular to it. These Feshbach resonances were found to be in excellent agreement with the complex poles of the adiabatic one-dimensional potential as described here in Eq.2. The narrowest resonance is located at the top of the DB potential and describe the TS. Beside the physical insight we get by the use of the adiabatic approach we also gain in the drastic reduction of the computational effort which is required for calculating the poles. Here we have shown that the non-resonance poles of 1D Eckart-type potential barrier (associated with complex scaled eigenfunctions which are localized in the forbidden classical region in 2D phase space) are the resonance poles of the 2D problem and are associated with complex-scaled eigenfunctions which are localized in the classically accessible region in the four-dimensional phase space. These type of resonances are marked here as DB resonances. The idea is that the Fabrey-Perot cavity couples the TS resonance with one of the DB resonances which have a broader width then the width of the TS and thereby the rate of the reaction $Ar+HCL\to H+Ar+Cl$ is enhanced.
@@ VI Possible Feasible Experiments for Enhancing the Rate of Gas-Phase Reactions in a Dark Cavity
The distance between the two mirrors is verified from a large distance where the interaction of the molecule with the cavity is practically zero, to an optimal distance where the standing QED wave has the same wavelength as the wavelength that couples the transition state (TS) complex pole and the dissociation barrier (DB) complex pole (which is inactive when the two mirrors are far apart).
@@
Condition (1b) is satisfied, for example, in the reaction $OH+HO_{2}\leftrightarrow O_{2}+H_{2}O$ (triplet reaction). See Table in Ref.31 where the vibrational energies perpendicular to the reaction coordinate at the two almost degenerate transition points (TS1 and TS2) are given and also the normal vibrational modes of $OH+HO_{2}$ and of $O_{2}+H_{2}O$. In this example, two nearly degenerate adiabatic potential energy surfaces are illustrated, as shown in Fig. 4 of Ref.32. In this work We have shown that for asymmetric reaction of $O+D_{2}$ the rate of the reaction is subcutaneously enhanced by the dark cavity.
@@ VII Concluding remarks
where $n_{j}=0$ (ground vibrational states). $V_{r.c}(X)$ is the static potential barrier for the reaction coordinate trajectory, i.e., it is solutions of the many-body electronic Hamiltonian. This approach has been previously used to study the collisions of hydrogen atom with $ArCl$ Van der Waals molecule[20] and for several collision reactions, such as in Ref.31. Moreover, NHQM is used to describe the resonance states (complex poles of the scattering matrix) that are used as a basis set within the adiabatic approach.)
Conditions for enhancing reaction rates are derived and presented in Section VI. Using the first condition we show that for the collision of hydrogen with the van der-Walls molecule the rate of the reaction is enhanced significantly inside a tuned dark cavity. Notice that there are no photons in the cavity, in addition, the reactants are in their ground electronic and vibrational states, i.e., it is indeed a dark cavity. These conditions have been applied also to the collision reaction of oxygen (in a triplet state) with deuterium diatomic molecule. We demonstrate a dramatic enhancement of the reaction rate in a dark cavity, where the distance between the two mirrors in the Fabrey Perot cavity are tuned. That is, the QED state energy, $\hbar\omega_{cav}$, equals the energy difference between two coupled resonance states. Finally, based on our analysis for the symmetric hydrogen exchanged reaction in methane the rate of the reaction cannot enhanced by a dark cavity. Hence here we provides guidelines for enhancing the rates of reactions within a dark cavity.
@@ Acknowledgements.
Dr. Arie Landau from the Helen Diller Center for Quantum Mechanics at the Technion is acknowledged for carefully reading several versions of this manuscript and for his valuable comments and suggestions to simplify the representation of the conditions that should help identify which reaction rates can be enhanced in a dark cavity.

### v3

@@ IV.1 Theory - the cavity effect on the rate of a reaction
Notice that both $\Gamma_{polariton}(\epsilon_{cav})$ and $\Gamma_{min}^{polariton}(\epsilon_{cav})$ are also functions of $\omega_{cav}$ as one can see from Fig. 11 which presents the rate of reaction inside the cavity as a function of the strength of the coupling of the activated complex $[AB]^{\#}$ with the cavity for two different values of $\omega_{cav}.$ It shows how the decay rate of the $O(triplet)+D_{2}\to OD+D$ reaction is increased inside the cavity when the distance between the two mirrors is tuned to the resonance condition. The coupling between the cavity and the system, in the acceleration gauge representation is used here, where the dipole transition is given by, $d_{{TS,DB}}(\omega_{cav})=\langle E_{TS}|\partial V_{ad}/\partial X|E_{DB}\rangle$ and the coupling strength parameter (in atomic unites) is given by $\alpha_{cav}=\epsilon_{cav}/\omega_{cav}^{2}$[28]. For the resonance condition we use one might be able to select several different cavity frequencies. $\omega_{cav}$ which couples the TS (transition state) resonance with a dynamical barrier (DB) broad resonance provided its position is lower than the position of the TS resonance (such that the TS in the vacuum of the cavity mode have about the same energy as the DB resonance that in the first excited quantized field mode of the cavity), Notice that the minimal distances between the two mirrors is given by, $L_{Z}=\pi c/\omega_{cav}\equiv\lambda_{cav}/2$. Therefore, the minimal distance depends on $\omega_{cav}$.
As expected by the analysis given in Eq.9, the enhancement of the reaction rate by the cavity is reduced as the coupling of the molecular state with the external field is increased. This is a special case of the generality of Bethe’s proof that there are situations where the increasing of the strength of the coupling of the system with external field not necessarily increase the rate of the reaction. See Fig.11 that shows the Bethe effect in the calculations of the effect of the cavity on the decay rate of the chemical reaction. That is, the increasing of the coupling strength parameter does not result necessarily in increasing the rate. When the difference between the resonance widths is sufficiently large (i.e., the TS is coupled with a broad DB resonance) the increasing of the strength of the coupling between the cavity and the molecules results in the suppression of the rate. As one can see from the results presented in Fig.11 the maximal enhancement of the rate of the reaction is obtained for a relative weak coupling between the cavity and the molecules.
@@
Using the adiabatic resonance eigenvalues and eigenfunctions for the $O+D_{2}$ reaction we have calculated and introduced in the previous section we evaluate the effect of the cavity on the reaction rate using Eq.13.
Figure 11: The effect of the cavity on the reaction rate, $\Gamma_{polariton}$, of the TS polaritions of $O+D_{2}\xrightarrow[]{\Gamma_{polariton}(\epsilon_{cav})}OD+D$ reaction. Reaction rate in Hartree and cavity-coupling parameter, in atomic units. The cavity be tuned into resonance condition between the states marked in blue in Fig.10 by adjusting the distance between the two Fabry-Perot mirrors (see Eq. 6). Rate in Hartree and coupling in atomic units. Notice that outside of the cavity (i.e., $\epsilon_{cav}=0$) the reaction rate is very small compared to the setup when the cavity is present. In other words, in this case the cavity effect on increasing the reaction rate is enormous even when the strength of the coupling between a single molecule and the cavity is weak. See Eq. 6, which links $\epsilon_{cav}$ to the distance between the two mirrors of the Fabry-Perot dark cavity.
@@ V The influence of a dark cavity on additional chemical reactions
An example for the first case case is scattering of hydrogen atom from the Van der Waals complex ArCl[35]. The poles which are presented in Fig.2 of Ref.19 are resonances which are associated with the temporary trapping of a hydrogen atom in between Ar and Cl. Only the first row of resonances in this figure have a classical analogue and are associated with the classical unstable periodic orbits. All other resonances are quantum diffraction resonances which do not have a classical analogue. These are Feshbach resonances of the two-dimensional PES where one coordinate is the reaction coordinate and the second one is perpendicular to it. These Feshbach resonances were found to be in excellent agreement with the complex poles of the adiabatic one-dimensional potential as described here in Eq.2. The narrowest resonance is located at the top of the DB potential and describe the TS. Beside the physical insight we get by the use of the adiabatic approach we also gain in the drastic reduction of the computational effort which is required for calculating the poles. Here we have shown that the non-resonance poles of 1D Eckart-type potential barrier (associated with complex scaled eigenfunctions which are localized in the forbidden classical region in 2D phase space) are the resonance poles of the 2D problem and are associated with complex-scaled eigenfunctions which are localized in the classically accessible region in the four-dimensional phase space. These type of resonances are marked here as DB resonances. The idea is that the Fabry-Perot cavity couples the TS resonance with one of the DB resonances which have a broader width then the width of the TS and thereby the rate of the reaction $Ar+HCl\to H+Ar+Cl$ is enhanced.
@@ VI Possible Feasible Experiments for Enhancing the Rate of Gas-Phase Reactions in a Dark Cavity
The distance between the two mirrors is verified from a large distance where the interaction of the molecule with the cavity is practically zero, to an optimal distance where the standing quantized field mode has the same wavelength as the wavelength that couples the transition state (TS) complex pole and the dissociation barrier (DB) complex pole (which is inactive when the two mirrors are far apart).
@@
Condition (1b) is satisfied, for example, in the reaction $OH+HO_{2}\leftrightarrow O_{2}+H_{2}O$ (triplet reaction). See Table in Ref.36 where the vibrational energies perpendicular to the reaction coordinate at the two almost degenerate transition points (TS1 and TS2) are given and also the normal vibrational modes of $OH+HO_{2}$ and of $O_{2}+H_{2}O$. In this example, two nearly degenerate adiabatic potential energy surfaces are illustrated, as shown in Fig. 4 of Ref.37. In this work we have shown that for asymmetric reaction of $O+D_{2}$ the rate of the reaction is significantly increased by the dark cavity.
@@ VII Concluding remarks
where $n_{j}=0$ (ground vibrational states). $V_{r.c}(X)$ is the static potential barrier for the reaction coordinate trajectory, i.e., it is solutions of the many-body electronic Hamiltonian. This approach has been previously used to study the collisions of hydrogen atom with $ArCl$ Van der Waals molecule[19] and for several collision reactions, such as in Ref.36. Moreover, NHQM is used to describe the resonance states (complex poles of the scattering matrix) that are used as a basis set within the adiabatic approach.)
Conditions for enhancing reaction rates are derived and presented in Section VI. Using the first condition we show that for the collision of hydrogen with the van der-Walls molecule the rate of the reaction is enhanced significantly inside a tuned dark cavity. Notice that there are no photons in the cavity, in addition, the reactants are in their ground electronic and vibrational states, i.e., it is indeed a dark cavity. These conditions have been applied also to the collision reaction of oxygen (in a triplet state) with deuterium diatomic molecule. We demonstrate a dramatic enhancement of the reaction rate in a dark cavity, where the distance between the two mirrors in the Fabry Perot cavity are tuned. That is, the quantized field mode state energy, $\hbar\omega_{cav}$, equals the energy difference between two coupled resonance states. Finally, based on our analysis for the symmetric hydrogen exchanged reaction in methane the rate of the reaction cannot enhanced by a dark cavity. We have not presented details of this analysis, since the adiabatic 1D effective potential and the 1D potential obtained from electronic structure calculations along the 1D curvilinear reaction coordinate are almost identical, and, therefore, no effect of the cavity on the reaction rate can be observed. Hence here we provide general guidelines for enhancing the rates of reactions inside a dark cavity.
@@ Acknowledgements.
Dr. Arie Landau from the Helen Diller Center for Quantum Mechanics at the Technion is acknowledged for carefully reading several versions of this manuscript and for his valuable comments and suggestions to simplify the representation of the conditions that should help identify which reaction rates can be enhanced in a dark cavity. Dr. Suheir Assady from RAMBAM is acknowledged for her hospitality in her department during the major portion of this research. Besides providing me with privacy, she also gave me the peace of mind necessary to conduct my scientific work by applying to biology what is known in physics and chemistry as the von Neumann and Wigner Non-Crossing Rule. This rule suggests that when several unexpected phenomena are observed simultaneously, they are most likely correlated, and their correlation typically requires the optimization of at least two parameters.